Periodic Inventory Using FIFO, LIFO, and Weighted Average Cost Methods

The units of an item available for sale during the year were as follows:

Jan. 1 Inventory 14 units at $25 $350
Aug. 7 Purchase 15 units at $26 390
Dec. 11 Purchase 15 units at $28 420
44 units $1,160

There are 20 units of the item in the physical inventory at December 31. The periodic inventory system is used. Determine the inventory cost using (a) the first-in, first-out (FIFO) method; (b) the last-in, first-out (LIFO) method; and (c) the weighted average cost method (round per unit cost to two decimal places and your final answer to the nearest whole dollar).

a.First-in, first-out (FIFO):

b.Last-in, first-out (LIFO): 506

c.Weighted average cost:
Answer #1
Average cost per unit 26.36 =1160/44
a.First-in, first-out (FIFO) 550 =(15*28)+(20-15)*26
b.Last-in, first-out (LIFO) 506 =(14*25)+(20-14)*26
c.Weighted average cost 527 =20*26.36